OpenStudy (anonymous):

-21+b=-23 solve each equation

OpenStudy (anonymous):

seriously?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

-21+b=-23 -21+b+21=-23+21 b=-2